OpenStudy (anonymous):

Solve the inequality. 3/10 > k - 3/5 -

OpenStudy (anonymous):

the dash is under the > symbol

OpenStudy (anonymous):

k < 9/10 -

jhonyy9 (jhonyy9):

3/10 +3/5 >= k (3+6)/10 >=k 9/10 >=k

OpenStudy (anonymous):

the > sign means greater than OR equal to. It is important to recognize that -